OpenStudy (anonymous):

find the derivative of 2-5x using the limit definition. Which I believe to be (f(x+h)-f(x))/h.

OpenStudy (luffingsails):

Don't forget about the part where h->0 once you are finished. <laugh> The key christal is to do exactly what you have written for the equation you are given. Then to take the limit as h->0. f(x+h) = 2-5(x+h) f(x) = 2-5(x) Substitute, simplify and take the limit. Give it a try and see where it takes you.

OpenStudy (anonymous):

i tried and i got -5(2x+h)/h

OpenStudy (whpalmer4):

\[ f(x) = 2-5x, f(x+h) = 2-5(x+h) = 2 - 5x - 5h\] \[\lim_{h \rightarrow 0} ((2-5x-5h) - (2-5x))/h = \lim_{h \rightarrow 0} -5h/h = -5\]

OpenStudy (anonymous):

oh okk for some reason i didnt cancel out the 5x's. thanks!